Let D = {(x, y) ∈ R2: x > 0, 0 < y < x3}. Define

f(x, y) = {0, (x, y) ∈/ D

{1, (x, y) ∈ D.


a) Approaching (0, 0) along with the line y = mx for each real number m and the y-axis, prove that lim(x,y)→(0,0) f(x, y) exists and compute the limit.

b) Argue whether f is continuous at (0, 0).


1
Expert's answer
2021-09-27T18:17:09-0400

(a)

"\lim\limits_{(x,y)\rightarrow (0,0)} f(x,y)\\
=\lim\limits_{(x,y)\rightarrow (0,0)} f(x,mx)\\
So, \lim\limits_{(x,y)\rightarrow (0,0)} f(x,0)\leq\lim\limits_{(x,y)\rightarrow (0,0)} f(x,y)\leq\lim\limits_{(x,y)\rightarrow (0,0)} (mx)^3\\
\Rightarrow 0\leq\lim\limits_{(x,y)\rightarrow (0,0)} f(x,y)\leq0\\
Hence, \lim\limits_{(x,y)\rightarrow (0,0)} f(x,y)=0"

(b) Yes, the function is continuous at "x=0."